I am trying to find out how to find the slope of the line that contains
(9,5) and (9,9). I got this wrong on a text.
Thank you

Found 2 solutions by checkley75, venugopalramana:
Answer by checkley75(3666)

THE SLOPE IS
(Y2-Y1)/(X2-X1)
(9-5)/(9-9)
4/0 THIS INDICATES A VERTICLE LINE THROUGH THE POINT X=9. (IF YOU DIVIDED 9 BY 0 YOU'D GET INFINITY).
TO PROVE IT PLOT THE TWO POINTS AND DRAW A LINE THROUGH THEM.

Answer by venugopalramana(3286)

am trying to find out how to find the slope of the line that contains
(9,5)(X1,Y1) and (9,9)(X2,Y2). I got this wrong on a text.
Thank you
SLOPE = (Y2-Y1)/(X2-X1) = (9-5)/(9-9)=4/0= INFINITY....LINE IS PERPENDICULAR TO X AXIS OR PARALLEL TO Y AXIS.